# A fast and exact algorithm for stabilizer Rényi entropy via XOR–FWHT

### v1

@@ I Introduction
Quantum advantage refers to the possibility that a quantum computer achieves an exponential speedup over classical computation for specific tasks. This capability is commonly understood within the framework of resource theories [RevModPhys.91.025001] in which a set of key quantum resources collectively underpins computational advantage [RevModPhys.91.025001, shor, NielsenChuang2010, Daley2022, huang2025vastworldquantumadvantage, lanes2025frameworkquantumadvantage]. Among the resources most closely tied to quantum advantage are entanglement [Entanglement1, Entanglement2] and nonstabilizerness [chitambar2019quantum, liu2022manybody, PhysRevLett.128.050402, Leone_2024, tirrito2024quantifying, Turkeshi_2023, turkeshi2025pauli, Turkeshi_2024_2, jasser2025stabilizerentropyentanglementcomplexity, viscardi2025interplayentanglementstructuresstabilizer, Iannotti2025entanglement, cusumano2025nonstabilizernessviolationschshinequalities, bittel2025operationalinterpretationstabilizerentropy, varikuti2025impactcliffordoperationsnonstabilizing, tirrito2025anticoncentrationnonstabilizernessspreadingergodic, zhang2025stabilizerrenyientropytransition, qian2025quantum, moca2025nonstabilizernessgenerationmultiparticlequantum, dowling2025magic, bera2025nonstabilizernesssachdevyekitaevmodel, masotllima2024stabilizer, aditya2025mpembaeffectsquantumcomplexity, hernándezyanes2025nonstabilizernessquantumenhancedmetrologicalprotocols, falcão2025magicdynamicsmanybodylocalized, sticlet2025nonstabilizernessopenxxzspin, tirrito2025universalspreadingnonstabilizernessquantum, zhang2024quantummagicdynamicsrandom, cao2025gravitationalbackreactionmagical, Tarabunga_2024, qiant25, qian2025, hang2025, qiant2024a, frau2025, fan2025disentangling, huang2024cliffordcircuitsaugmentedmatrix, ding2025evaluating, korbany2025longrangenonstabilizernessphasesmatter, tarabunga2025efficientmutualmagicmagic, szombathy2025independentstabilizerrenyientropy, hou2025stabilizerentanglementenhancesmagic, hoshino2025stabilizerrenyientropyconformal, tarabunga2024magictransitionmeasurementonlycircuits, tirrito2025magicphasetransitionsmonitored, wang2025magictransitionmonitoredfree, santra2025complexitytransitionschaoticquantum, Haug2025probingquantum, Haug2023stabilizerentropies, Haug_2023_1, Lami_2023_2, Lami_2024, tarabunga2023manybody, tarabunga2024nonstabilizerness], which is also known as quantum magic. These resources are widely regarded as necessary ingredients for universal quantum computation [PhysRevLett.110.060504], and they play a central role in the construction and optimization of practical quantum devices [PhysRevLett.100.030504, PhysRevLett.91.147902, PhysRevLett.93.040502, PRXQuantum.3.020328]. It is important to emphasize that strong entanglement alone does not guarantee quantum advantage, since stabilizer states can be highly entangled while remaining efficiently simulable on classical computers under the Gottesman-Knill theorem [gottesman1998heisenbergrepresentationquantumcomputers, GT]. Nonstabilizerness, therefore, serves as an independent resource that quantifies the departure from the stabilizer manifold and that reflects the consumption of non-Clifford resources required to prepare a given state, which is why it is viewed as a particularly scarce and crucial resource for universal quantum computation when non-Clifford operations are costly in experiments and in fault-tolerant architectures [Preskill2018NISQ, PRXQuantum.3.020333, Quantinuum2025, Aasen2025, Peham2025].
@@
In recent years, the behavior of magic and related nonstabilizerness quantifiers in many-body quantum systems has also become an active topic at the interface of quantum information [HaydenPreskill2007JHEP, SekinoSusskind2008JHEP, HosurQiRobertsYoshida2016JHEP, vonKeyserlingkRakovszkyPollmannSondhi2018PRX, KhemaniVishwanathHuse2018PRX, Swingle2018NatPhys, LiuZhangYinZhang2024PRL, LiuZhangYinZhangYao2024arXiv, YuLiZhang2025CPL, HuangLiHuseChan2023Scholarpedia, PhysRevResearch.7.L012011, Chen2025subsystem, PhysRevB.110.064323, PhysRevLett.132.240402] and many-body physics [PhysRevA.106.042426, Tarabunga2024criticalbehaviorsof, PhysRevA.108.042407, PhysRevA.110.022436, Passarelli2025chaosmagicin, PhysRevB.110.045101, delafuente2025topologicalstabilizermodelscontinuous, PRXQuantum.4.040317, PhysRevB.111.L081102, 10.21468/SciPostPhys.18.5.165, PaviglianitiLamiColluraSilva2025PRXQ, PRXQuantum.5.030332, li2025nonstabilizernessstarkmanybodylocalization, xfp5-hhs4, 1jzy-sk9r, LarkinOvchinnikov1969JETP, ShenkerStanford2014JHEP, MaldacenaShenkerStanford2016JHEP, Swingle2018NatPhys, HashimotoMurataYoshii2017JHEP, RobertsYoshida2017JHEP, CotlerHunterJonesLiuYoshida2017JHEP, LeoneOlivieroZhouHamma2021Quantum, GotoNosakaNozaki2022PRD, PaviglianitiLamiColluraSilva2025PRXQ, Prosen2021Chaos, HuangLiHuseChan2023Scholarpedia, aditya2025growthspreadingquantumresources, Li_2025, lhz3, lhz2, PhysRevB.111.184302, zhao2025entanglementphasesphasetransitions, PhysRevA.109.062226, li2025quantummpembaeffectlongranged, sxz1, sxz2, sxz3, sxz4, liu2025noisymonitoredquantumcircuits, 6722-tf9c, yu2025gaplesssymmetryprotectedtopologicalstates, Koh_2022, Koh_2024, Tianqi, shenruizhe, digital1, digital2, digital3]. Quantifying nonstabilizerness is challenging both conceptually and computationally. A broad range of magic measures has been proposed, which includes quantifiers based on quasiprobability representations, robustness-type measures, and measures related to operational or simulation overhead [Veitch2014, WOOTTERS19871, Veitch_2012, PhysRevLett116250501, Bravyi, PhysRevLett.124.090505, Heimendahl2021stabilizerextentis]. For generic many-body states, however, evaluating these quantifiers typically entails exponential computational cost, which severely limits their applicability to larger system sizes.
@@
$\mathcal{M}_{2}(|\psi\rangle)=-\frac{1}{d}\log\!\left(\sum_{P\in\mathcal{P}_{N}}\bigl|\langle\psi|P|\psi\rangle\bigr|^{4}\right),$ (1)
@@
The structure of this work is organized as follows. In Sec. II, starting from Eq. (1), we present a detailed derivation. After representing Pauli strings as bitstrings, we systematically simplify the expression by means of XOR convolution and the WHT, ultimately arriving at a compact form that is well suited for numerical implementation and exhibits significantly improved computational efficiency. In Sec. III, we provide the pseudocode of the proposed XOR-FWHT algorithm, and present the corresponding Python implementation details in the Appendix, together with benchmarks of the actual runtime performance on multicore CPU architectures. In Sec. IV, we present a series of numerical results, which verify the consistency of our algorithm with both theoretical predictions and brute-force approaches; moreover, the high numerical precision of the method enables it to resolve subtle effects arising from small parameter variations in physical models. Finally, in Sec. V, we summarize the main results of this work and discuss potential directions for future developments of the XOR-FWHT algorithm.
@@ III XOR-FWHT Algorithm for the $2$-SREs
We benchmarked the performance of the algorithm on a workstation equipped with two Intel Xeon Platinum 8352V processors (36 cores per CPU, 72 cores in total). In the benchmarks, the input states were randomly generated Haar states obtained by normalizing complex Gaussian random vectors, and we measured the wall-clock time for a single evaluation of the 2-SRE. For different system sizes $N$, the typical runtimes are as follows: approximately $0.88$ seconds for $N=14$; about $12$ seconds for $N=16$; about $300$ seconds for $N=18$; and about $7200$ seconds for $N=20$. These numerical results are consistent with the theoretical time complexity $O(N4^{N})$ of the algorithm.
@@
IV Numerical Results and Physical Examples
In this section, we present several representative numerical results. First, we randomly sample Haar-random pure states from the Hilbert space according to the Haar measure. As typical quantum states, Haar-random states exhibit volume-law entanglement near the Page limit at the subsystem level, which causes tensor-network-based approaches to compute $\mathcal{M}_{2}$ to encounter an exponentially growing bond-dimension bottleneck as the system size increases. In contrast, the XOR-FWHT algorithm proposed in this work does not rely on the bond dimension of a tensor-network representation, but instead acts directly on the state vector itself, and therefore enables efficient evaluation of the magic measure $\mathcal{M}_{2}$ for Haar-random states. As shown in Fig. 1(a), the numerical results are in good agreement with the theoretical prediction [PhysRevB.111.054301], thereby validating the correctness and applicability of the algorithm for highly entangled states.
Figure 1: (a) The average 2-SRE $\mathcal{M}_{2}$ of Haar-random pure states as a function of the system size $N$. For each NN, we independently sample 1010010-100 Haar-random states and compute the average. The solid line shows the theoretically predicted behavior $\log_{2}(2^{N}+3)-2$, which is in good agreement with the numerical results, indicating that $\mathcal{M}_{2}$ of Haar states grows approximately linearly with NN. (b) Time evolution of $\mathcal{M}_{2}$ in the dynamics of the XXZ model, comparing results obtained from the brute-force approach and from the XOR-FWHT algorithm proposed in this work. The two methods yield identical numerical results throughout the entire time evolution, verifying the correctness and stability of the XOR-FWHT algorithm in dynamical calculations.
As shown in Fig. 1(b), we consider a one-dimensional XXZ spin-$\frac{1}{2}$ chain with the Hamiltonian
@@ IV Numerical Results and Physical Examples
We next consider the transverse-field Ising model with an additional longitudinal field (TFIM with longitudinal field), whose Hamiltonian is given by
$H_{\mathrm{TFIM-LF}}=-J\sum_{i=0}^{N-1}Z_{i}Z_{i+1}-h_{x}\sum_{i=0}^{N-1}X_{i}-h_{z}\sum_{i=0}^{N-1}Z_{i},$ (24)
@@ A.1 Definition
A.2 Property: Involution and Inverse Transform
@@ A.5 Property II: Parseval’s Theorem
The WHT preserves the $\ell_{2}$ norm (up to a fixed constant factor):

### v2

# A fast and exact approach for stabilizer Rényi entropy via the XOR–FWHT algorithm

@@ I Introduction
Quantum advantage refers to the possibility that a quantum computer achieves an exponential speedup over classical computation for specific tasks. This capability is commonly understood within the framework of resource theories [RevModPhys.91.025001] in which a set of key quantum resources collectively underpins computational advantage [RevModPhys.91.025001, shor, NielsenChuang2010, Daley2022, huang2025vastworldquantumadvantage, lanes2025frameworkquantumadvantage]. Among the resources most closely tied to quantum advantage are entanglement [Entanglement1, Entanglement2] and nonstabilizerness [chitambar2019quantum, liu2022manybody, PhysRevLett.128.050402, Leone_2024, tirrito2024quantifying, Turkeshi_2023, turkeshi2025pauli, Turkeshi_2024_2, jasser2025stabilizerentropyentanglementcomplexity, viscardi2025interplayentanglementstructuresstabilizer, Iannotti2025entanglement, cusumano2025nonstabilizernessviolationschshinequalities, bittel2025operationalinterpretationstabilizerentropy, varikuti2025impactcliffordoperationsnonstabilizing, tirrito2025anticoncentrationnonstabilizernessspreadingergodic, zhang2025stabilizerrenyientropytransition, qian2025quantum, moca2025nonstabilizernessgenerationmultiparticlequantum, dowling2025magic, bera2025nonstabilizernesssachdevyekitaevmodel, masotllima2024stabilizer, aditya2025mpembaeffectsquantumcomplexity, hernándezyanes2025nonstabilizernessquantumenhancedmetrologicalprotocols, falcão2025magicdynamicsmanybodylocalized, sticlet2025nonstabilizernessopenxxzspin, tirrito2025universalspreadingnonstabilizernessquantum, zhang2024quantummagicdynamicsrandom, cao2025gravitationalbackreactionmagical, Tarabunga_2024, qiant25, qian2025, hang2025, qiant2024a, frau2025, fan2025disentangling, huang2024cliffordcircuitsaugmentedmatrix, ding2025evaluating, korbany2025longrangenonstabilizernessphasesmatter, wei2025longrangenonstabilizernessquantumcodes, tarabunga2025efficientmutualmagicmagic, szombathy2025independentstabilizerrenyientropy, hou2025stabilizerentanglementenhancesmagic, hoshino2025stabilizerrenyientropyconformal, tarabunga2024magictransitionmeasurementonlycircuits, tirrito2025magicphasetransitionsmonitored, wang2025magictransitionmonitoredfree, santra2025complexitytransitionschaoticquantum, Haug2025probingquantum, Haug2023stabilizerentropies, Haug_2023_1, Lami_2023_2, Lami_2024, tarabunga2023manybody, tarabunga2024nonstabilizerness], which is also known as quantum magic. These resources are widely regarded as necessary ingredients for universal quantum computation [PhysRevLett.110.060504], and they play a central role in the construction and optimization of practical quantum devices [PhysRevLett.100.030504, PhysRevLett.91.147902, PhysRevLett.93.040502, PRXQuantum.3.020328]. It is important to emphasize that strong entanglement alone does not guarantee quantum advantage, since stabilizer states can be highly entangled while remaining efficiently simulable on classical computers under the Gottesman-Knill theorem [gottesman1998heisenbergrepresentationquantumcomputers, GT]. Nonstabilizerness, therefore, serves as an independent resource that quantifies the departure from the stabilizer manifold and that reflects the consumption of non-Clifford resources required to prepare a given state, which is why it is viewed as a particularly scarce and crucial resource for universal quantum computation when non-Clifford operations are costly in experiments and in fault-tolerant architectures [Preskill2018NISQ, PRXQuantum.3.020333, Quantinuum2025, Aasen2025, Peham2025].
@@
In recent years, the behavior of magic and related nonstabilizerness quantifiers in many-body quantum systems has also become an active topic at the interface of quantum information [HaydenPreskill2007JHEP, SekinoSusskind2008JHEP, HosurQiRobertsYoshida2016JHEP, vonKeyserlingkRakovszkyPollmannSondhi2018PRX, KhemaniVishwanathHuse2018PRX, Swingle2018NatPhys, LiuZhangYinZhang2024PRL, LiuZhangYinZhangYao2024arXiv, YuLiZhang2025CPL, HuangLiHuseChan2023Scholarpedia, PhysRevResearch.7.L012011, Chen2025subsystem, PhysRevB.110.064323, PhysRevLett.132.240402] and many-body physics [PhysRevA.106.042426, Tarabunga2024criticalbehaviorsof, PhysRevA.108.042407, PhysRevA.110.022436, Passarelli2025chaosmagicin, PhysRevB.110.045101, delafuente2025topologicalstabilizermodelscontinuous, PRXQuantum.4.040317, PhysRevB.111.L081102, 10.21468/SciPostPhys.18.5.165, PaviglianitiLamiColluraSilva2025PRXQ, PRXQuantum.5.030332, li2025nonstabilizernessstarkmanybodylocalization, xfp5-hhs4, 1jzy-sk9r, LarkinOvchinnikov1969JETP, ShenkerStanford2014JHEP, MaldacenaShenkerStanford2016JHEP, Swingle2018NatPhys, HashimotoMurataYoshii2017JHEP, RobertsYoshida2017JHEP, CotlerHunterJonesLiuYoshida2017JHEP, LeoneOlivieroZhouHamma2021Quantum, GotoNosakaNozaki2022PRD, PaviglianitiLamiColluraSilva2025PRXQ, Prosen2021Chaos, HuangLiHuseChan2023Scholarpedia, maity2025localspreadingstabilizerrenyi, aditya2025growthspreadingquantumresources, Li_2025, lhz3, lhz2, PhysRevB.111.184302, zhao2025entanglementphasesphasetransitions, PhysRevA.109.062226, li2025quantummpembaeffectlongranged, sxz1, sxz2, sxz3, sxz4, liu2025noisymonitoredquantumcircuits, 6722-tf9c, yu2025gaplesssymmetryprotectedtopologicalstates, Koh_2022, Koh_2024, Tianqi, shenruizhe, digital1, digital2, digital3]. Quantifying nonstabilizerness is challenging both conceptually and computationally. A broad range of magic measures has been proposed, which includes quantifiers based on quasiprobability representations, robustness-type measures, and measures related to operational or simulation overhead [Veitch2014, WOOTTERS19871, Veitch_2012, PhysRevLett116250501, Bravyi, PhysRevLett.124.090505, Heimendahl2021stabilizerextentis]. For generic many-body states, however, evaluating these quantifiers typically entails exponential computational cost, which severely limits their applicability to larger system sizes.
@@
$\mathcal{M}_{2}(|\psi\rangle)=-\log\!\left(\frac{1}{d}\ \sum_{P\in\mathcal{P}_{N}}\bigl|\langle\psi|P|\psi\rangle\bigr|^{4}\right),$ (1)
@@
The structure of this work is organized as follows. In Sec. II, starting from Eq. (1), we present a detailed derivation. After representing Pauli strings as bitstrings, we systematically simplify the expression by means of XOR convolution and the WHT, ultimately arriving at a compact form that is well suited for numerical implementation and exhibits significantly improved computational efficiency. In Sec. III, we provide the pseudocode of the proposed XOR-FWHT algorithm, and present the corresponding Python implementation details in the Appendix, together with benchmarks of the actual runtime performance on multicore CPU architectures. In Sec. IV, we present a series of numerical applications, which verify the consistency of our algorithm with both theoretical predictions and brute-force approaches; moreover, the high numerical precision of the method enables it to resolve subtle effects arising from small parameter variations in physical models. Finally, in Sec. V, we summarize the main results of this work and discuss potential directions for future developments of the XOR-FWHT algorithm.
@@ III XOR-FWHT Algorithm for the $2$-SREs
We benchmarked the performance of the algorithm on a workstation equipped with two Intel Xeon Platinum 8352V processors (36 cores per CPU, 72 cores in total). In the benchmarks, the input states were randomly generated Haar states obtained by normalizing complex Gaussian random vectors, and we measured the wall-clock time for a single evaluation of the 2-SRE. For different system sizes $N$, the typical runtimes are as follows: approximately $0.88$ seconds for $N=14$; about $12$ seconds for $N=16$; about $300$ seconds for $N=18$; and about $2$ hours for $N=20$. These numerical results are consistent with the theoretical time complexity $O(N4^{N})$ of the algorithm.
@@
IV Applications in Physical Examples
In this section, we numerically present several representative physical examples by using algorithm 1 for evaluating 2-SRE.
IV.1 Example 1. validation on Haar random states
First, we randomly sample Haar-random pure states from the Hilbert space according to the Haar measure. As typical quantum states, Haar-random states exhibit volume-law entanglement near the Page limit at the subsystem level, which causes tensor-network-based approaches to compute $\mathcal{M}_{2}$ to encounter an exponentially growing bond-dimension bottleneck as the system size increases. In contrast, the XOR-FWHT algorithm proposed in this work does not rely on the bond dimension of a tensor-network representation, but instead acts directly on the state vector itself, and therefore enables efficient evaluation of the magic measure $\mathcal{M}_{2}$ for Haar-random states.
As shown in Fig. 1(a), the numerical results are in good agreement with the theoretical prediction [PhysRevB.111.054301], thereby validating the correctness and applicability of the algorithm for highly entangled states.
Figure 1: (a) The average 2-SRE $\mathcal{M}_{2}$ of Haar-random pure states as a function of the system size $N$. For each NN, we independently sample 1010010-100 Haar-random states and compute the average. The solid line shows the theoretically predicted behavior $\log_{2}(2^{N}+3)-2$, which is in good agreement with the numerical results, indicating that $\mathcal{M}_{2}$ of Haar states grows approximately linearly with $N$. (b) Time evolution of $\mathcal{M}_{2}$ in the dynamics of the XXZ model, comparing results obtained from the brute-force approach and from the XOR-FWHT algorithm proposed in this work. The two methods yield identical numerical results throughout the entire time evolution, verifying the correctness and stability of the XOR-FWHT algorithm in dynamical calculations.
IV.2 Example 2. Hamiltonian quench dynamics
Next, we consider a one-dimensional XXZ spin-$\frac{1}{2}$ chain with the Hamiltonian
@@ IV Applications in Physical Examples
We also consider the transverse-field Ising model with an additional longitudinal field (TFIM with longitudinal field), whose Hamiltonian is given by
$H_{\mathrm{TFIM+LF}}=-J\sum_{i=0}^{N-1}Z_{i}Z_{i+1}-h_{x}\sum_{i=0}^{N-1}X_{i}-h_{z}\sum_{i=0}^{N-1}Z_{i},$ (24)
@@
IV.3 Example 3. random unitary circuit quench dynamics
[…]
As illustrated in Fig. 3b, we numerically simulated this setup using TensorCircuit-NG [Zhang_2023] for system sizes $N=10,12,14,16$. For all system sizes considered, the SRE dynamics saturates at long times to the corresponding Haar value.
@@ A.1 Definition
A.2 Property of Involution and Inverse Transform
@@ A.5 Property II: Parseval’s Theorem
The WHT preserves the $\ell_{2}$ norm: